# Insights into the Incorporation of Signal Information in Binaural Signal Matching with Wearable Microphone Arrays

### v2

@@ VI-B Methodology
The binaural signals in this experiment were generated according to Section V-B. The study employed the Multiple Stimuli with Hidden Reference and Anchor (MUSHRA) test [27]. The test featured two MUSHRA screens: both with $50^{\circ}$ head rotation, one with an error of $10^{\circ}$ in the estimation of speaker direction, as shown in Section V-E and one without. Within each screen, the reference signal was a HOA signal of order $N=14$, detailed in Section V-B. Each screen included the following four test signals: the hidden reference (HOA), the anchor BSM, d-BSM, and COM, implemented as detailed in Section V-B. Ten subjects, all reported to have normal hearing, participated in the study and were presented with MUSHRA screens and signals in a randomized sequence. They evaluated the similarity of each test signal to the reference signal based on overall quality, encompassing spatial and timbral characteristics. Scores ranged from 0 to 100, with 100 indicating that the test signal was indistinguishable from the reference. In addition, headphone compensation filters cited in [18] were utilized. Prior to the listening test, participants underwent two training stages: the first familiarized them with the scoring criteria, while the second introduced them to the test signals without requiring scoring. A Repeated Measures ANalysis of VAriance (RM-ANOVA) [42] with two within-subject factors and their interaction was conducted on the scores.
@@ VII Conclusions
This paper offers several significant contributions to the field of binaural signal matching and its application with wearable microphone arrays. First, a new performance metric for BSM, termed the “directional error,” is introduced. This metric quantifies the expected relative error as a function of sound source direction, revealing that BSM errors are highly dependent on the DOA of sound sources, which is closely related to the HRTF magnitude. This behavior was further validated using ITD and ILD analysis. Additionally, the paper presents, analyzes, and compares two approaches that incorporate signal information into the BSM method, specifically focusing on source DOA. These methods effectively address the limitations imposed by the directional behavior of BSM, demonstrating improved performance across all source directions. This enhanced performance was also confirmed through ITD and ILD analysis. Furthermore, the performance analysis revealed a trade-off where enhancing BSM performance for the DOA of a sound source leads to a slight decrease in performance across other directions. One of the proposed methods showed superior performance in directions away from the source and was found to be more robust to errors in the estimation of DOA. Finally, the paper includes a comprehensive simulation study and a listening experiment that validate the accuracy and quality of the different methods using a wearable microphone array. The results confirm the effectiveness of the proposed approaches and highlight their potential for practical applications. Despite these advances, several limitations remain. The performance of the proposed methods can be significantly affected by errors in DOA estimation, which was evident from the robustness analysis. Additionally, the methods were primarily tested in controlled environments, and their performance in more complex real-world scenarios remains to be fully explored. Future work could focus on improving DOA estimation accuracy to further enhance the robustness of the proposed methods. Additionally, extending the validation of these methods to diverse and dynamic real-world environments would be valuable. Investigating the integration of these methods with other spatial audio processing techniques could also offer new insights and improvements in binaural signal reproduction.

### v3

@@ VI-B Methodology
The binaural signals in this experiment were generated according to Section V-B. The study employed the Multiple Stimuli with Hidden Reference and Anchor (MUSHRA) test [27]. The test featured two MUSHRA screens: both with $50^{\circ}$ head rotation, one with an error of $10^{\circ}$ in the estimation of speaker direction, as shown in Section V-D and one without. Within each screen, the reference signal was a HOA signal of order $N=14$, detailed in Section V-B. Each screen included the following four test signals: the hidden reference (HOA), the anchor BSM, d-BSM, and COM, implemented as detailed in Section V-B. Ten subjects, all reported to have normal hearing, participated in the study and were presented with MUSHRA screens and signals in a randomized sequence. They evaluated the similarity of each test signal to the reference signal based on overall quality, encompassing spatial and timbral characteristics. Scores ranged from 0 to 100, with 100 indicating that the test signal was indistinguishable from the reference. In addition, headphone compensation filters cited in [18] were utilized. Prior to the listening test, participants underwent two training stages: the first familiarized them with the scoring criteria, while the second introduced them to the test signals without requiring scoring. A Repeated Measures ANalysis of VAriance (RM-ANOVA) [42] with two within-subject factors and their interaction was conducted on the scores.
@@ VII Conclusions
This paper offers several significant contributions to the field of binaural signal matching and its application with wearable microphone arrays. The paper presents, analyzes, and compares two approaches that incorporate signal information into the BSM method, specifically focusing on source DOA. These methods effectively address the limitations imposed by the diffuse-field assumption of BSM, demonstrating improved performance across source directions. This enhanced performance was also confirmed through ITD and ILD analysis. Furthermore, the performance analysis revealed a only a minor trade-off, where enhancing BSM performance for the DOA of a sound source leads to a slight decrease in performance across other directions. One of the proposed methods showed superior performance in directions away from the source and was found to be more robust to errors in the estimation of DOA. Finally, the paper includes a comprehensive simulation study and a listening experiment that validate the accuracy and quality of the different methods using a wearable microphone array. The results confirm the effectiveness of the proposed approaches and highlight their potential for practical applications. Despite these advances, several limitations remain. The performance of the proposed methods can be significantly affected by errors in DOA estimation, which was evident from the robustness analysis. Additionally, the methods were primarily tested in controlled environments, and their performance in more complex real-world scenarios remains to be fully explored. Future work could focus on improving DOA estimation accuracy to further enhance the robustness of the proposed methods. Additionally, extending the validation of these methods to diverse and dynamic real-world environments would be valuable. Investigating the integration of these methods with other spatial audio processing techniques could also offer new insights and improvements in binaural signal reproduction.